### v2

# AI Hospital: Interactive Evaluation and Collaboration of LLMs as Intern Doctors for Clinical Diagnosis

Abstract
The incorporation of Large Language Models (LLMs) in healthcare marks a significant advancement. However, the application has predominantly been limited to discriminative and question-answering tasks, which does not fully leverage their interactive potential. To address this limitation, our paper presents AI Hospital, a framework designed to build a real-time interactive diagnosis environment. To simulate the procedure, we collect high-quality medical records to create patient, examiner, and medical director agents. AI Hospital is then utilized for the interactive evaluation and collaboration of LLMs. Initially, we create a Multi-View Medical Evaluation (MVME) benchmark where various LLMs serve as intern doctors for interactive diagnosis. Subsequently, to improve diagnostic accuracy, we introduce a collaborative mechanism that involves iterative discussions and a dispute resolution process under the supervision of the medical director. In our experiments, we validate the reliability of AI Hospital. The results not only explore the feasibility of apply LLMs in clinical consultation but also confirm the effectiveness of the dispute resolution focused collaboration method. [^1] https://github.com/LibertFan/AI_Hospital
1 Introduction
The realm of healthcare has witnessed a remarkable evolution with the integration of artificial intelligence Bajwa et al. (2021), particularly through the development of Large Language Models (LLMs) OpenAI (2023); Touvron et al. (2023); Bai et al. (2023). These advancements open novel pathways for medicine Singhal et al. (2023); Chen et al. (2023e); Bao et al. (2023); Zhang et al. (2023), with LLMs demonstrating remarkable performance across various medical question-answering benchmarks Jin et al. (2021, 2019); Pal et al. (2022). Nonetheless, the complexity of real-world medical consultations, which typically involve a series of steps such as assessing patient conditions, performing physical and auxiliary examinations, and offering medication and health guidance, presents a challenge. These steps require intricate, multi-turn interactions with patients, raising questions about the efficacy of LLMs in managing real-world medical consultations.
In this work, we introduce AI Hospital, a platform based on multi-agent interaction Park et al. (2023); Li et al. (2023) to build a real-time interactive consultation scenario, and explore the feasibility of using LLMs as doctors in clinical diagnosis. The platform incorporates various roles including intern doctors, patients, examiners, and a medical director. In the framework, intern doctors diagnose diseases through active questioning, patients need to describe their symptoms and respond to inquiries from doctors, examiners execute medical tests following the doctors’ instructions, while the medical director is tasked with assessing the performance of the intern doctors in diagnosis. Significantly, the primary objective of AI Hospital is twofold: firstly, facilitating interactive consultation among intern doctors, patients and examiners to evaluate the diagnostic capability of LLMs, and secondly, enabling collaborative efforts among doctors, thereby improving the diagnostic accuracy.
Figure 1: The demonstration of AI Hospital.
Establishing AI Hospital necessitates overcoming the challenge of replicating the behavior of personnel in medical settings, such as patients and examiners. (1) To address the issue of medical roles lacking medical knowledge, we collect high-quality Chinese medical records, utilizing professional medical data as inputs for large language models to effectively simulate the agents. (2) Given the critical role of professional medical examinations in clinical diagnosis, with CDC finding that 70% medical decisions rely on them di Ruffano et al. (2012); CDC (2018), we establish an examiner role specifically for recording and overseeing patients’ physical and auxiliary examination results. This configuration also restricts patient agents from accessing detailed medical examination data, ensuring the alignment in behavior with actual patients. (3) To mitigate patients’ lack of self-awareness in conversation scenarios Wang et al. (2023c), we craft background narratives, personas and descriptions of conditions with personal characteristics to guide the conversational style. In summary, we intricately craft the prompts for these medical roles, allowing for natural interaction with intern doctors grounding on the setting of role.
Based on the AI Hospital, we investigate the feasibility of utilizing LLMs as intern doctors for clinical diagnosis by establishing a Multi-View Medical Evaluation (MVME) benchmark. The benchmark assesses the real-time interactive consultation capabilities of LLMs across three critical dimensions. Firstly, it examines their ability to identify patient symptoms, highlighting the importance of actively seeking relevant information. Secondly, we assess the comprehensiveness of their medical examinations, specifically their adeptness in selecting and administering a suitable range of these tests. Thirdly, we measure the accuracy and professionalism of their diagnosis, checking whether they meet standards of medical practice. In practice, we not only leverage GPT-4 as a flexible evaluation engine but also employ a link-based method to incorporate medical standard knowledge.
To enhance LLMs’ diagnostic accuracy, drawing on previous research Croft et al. (2015); Llewelyn et al. (2014); Centor et al. (2019) that highlights the importance of teamwork in clinical diagnosis, we delve into the collaborative mechanism O’leary et al. (2010); Lamb et al. (2011), facilitating iterative discussions among intern doctors. Different from previous models that lack concise organization and clear direction for participants Talebirad and Nadiri (2023); Liang et al. (2023), our framework introduces a dispute resolution strategy that effectively engages the medical director to guide discussions, clarify issues, and steer the interns towards a more structured and efficient collaboration, thereby fostering more focused discussions and accelerating the achievement of consensus.
In our experiment, we initially investigate the behavior of various personnel within the AI Hospital to validate the reliability. Subsequently, we test a range of LLMs in the AI Hospital, uncovering that deficiencies in interaction significantly impair diagnostic accuracy. Furthermore, we assess the effectiveness of our dispute resolution-focused cooperative mechanism, demonstrating its capacity to accelerate and enhance diagnostic accuracy. Finally, we dive into GPT-4, analyzing its limited efficacy in identifying necessary medical examinations and managing complications.
In general, our contributions are of three-folds:
  • We develop AI Hospital underpinned by an LLM-driven multi-agent framework.
  • We build a benchmark MVME for interactive evaluation within the medical domain to assess LLMs in interactive consultations.
  • We design a collaborative mechanism which focuses on dispute resolution. It demonstrates diagnostic capabilities surpassing GPT-4.
2 Setup of AI Hospital
The framework of our AI Hospital is illustrated in Figure 1. It encompasses three principal procedures: (1) Intern doctors proactively question patients about their physical condition and patients obtain diagnostic reports from intern doctors; (2) Patients undertake medical examinations following the directions of intern doctors; (3) Medical director assess the performance of intern doctors. In this section, §2.1 details how we utilize the collected medical records to equip medical roles with necessary medical expertise. Subsequently, §2.2 to §2.5 describe the methodologies employed to establish the four agents, including the intern doctor.
2.1 Agents Setup on Medical Records
We collect a substantial volume of Chinese medical record data, comprising various sections. Given the requirement of specific medical knowledge in different roles, we categorize the information of medical record into three types.
Basic Information. It is consisting of chief complaints, current and past medical histories, and personal and familial histories.
Professional Medical Examination. It encompasses physical and auxiliary examinations.
Diagnosis and Treatment. It has diagnostic results, diagnostic rationales, and treatment course.
These categories are correspondingly allocated to the patient, examiner, and medical director, then employed in prompts of LLMs to facilitate the medical knowledge of distinct roles.
2.2 Intern Doctor
In AI Hospital, intern doctors are responsible for clinical diagnosis. This requires them to actively engage with patients, delving into detailed inquiries about their physical conditions. Through these interactions, intern doctors are expected to apply their medical knowledge and skills to identify potential illnesses afflicting the patients. Subsequently, the doctors conduct medical examinations through examiners to carry out a thorough investigation. Once a diagnosis is identified, the intern doctors then move to the formulation of treatment plans. They are expected to consider various treatment options, weigh their effectiveness, and select the most appropriate course of action. The prompt of intern doctor is shown in Table 12.
Figure 2: The workflow of examiner.
2.3 Examiner
In an effort to replicate the functions of a hospital’s examination department, we introduce the examiner agent for the administration of medical examination. It is designed to manage “Professional Medical Examination” as detailed in § 2.1. When patient agents ask for specific tests under physician guidance, the examiner agent should provide the relevant examination results.
To enhance the rationality of the inspection process and ensure that results are provided only upon receipt of correct examination items, we divide the examination procedure into two sequential stages as shown in Figure 2. In the first stage, examiners are tasked with accurately interpreting patients’ examination requests and discerning the required tests as the prompt in Table 10. This involves extracting essential examination items and dismissing requests with vague objectives, such as “inquiries about general medical examination conditions”. In the second phase, examiners follow the prompt in Table 11 to convey the examination outcomes to the doctors, contingent upon the examination items identified in the initial phase.
2.4 Patient
Functionality. We conduct a systematic examination of the essential capabilities required from patients during interactive diagnosis. These capabilities include: (1) the articulate expression of complaints; (2) truthful responses to doctors’ inquiries; (3) active participation in examinations; and (4) the voicing of questions and concerns. We tailor prompts accordingly to meet the requirements. The prompt is shown in Table 9.
Personality. To facilitate improved behavior control, we propose enhancing self-awareness in patient role-playing through the development of a patient’s personality. First, we utilize GPT-4 to reason and imagine the detailed aspects of the patient’s lifestyle, pertinent events, personality traits, and speaking manner that correlate with their physical condition. There is an example: “You are a farmer who left school after primary education. Your life is frugal, but your mindset is optimistic. You lack knowledge of medical terminology.” Second, in line with the patient’s profile, we adopt GPT-4 to generate a statement reflecting their condition. For example, “Doctor, I’ve been having this pain on the right side of my stomach, and I don’t want to eat at all. Also, I’ve been peeing blood these past couple of days.” This statement serves as the first segment of the interactive consultation for one-shot demonstration. By crafting comprehensive patient profiles as prompts and employing one-shot learning, the patient agent can replicate the behaviors of a real patient in dialogues.
2.5 Medical Director
The primary responsibility of the medical director is to assess the performance of intern doctors in interactive diagnosis. At the end of a consultation, the medical director asks the intern doctors to provide a comprehensive summary that includes the patient’s symptom, medical examinations, diagnostic results, diagnostic rationales, and the treatment plan. This procedure is designed to consolidate information regarding the patient’s state, elucidate the interns’ comprehension of the diagnosed condition, the basis of their judgments, and the treatment strategy. Upon receiving responses about these five areas, the medical director evaluates each section by comparing it with the medical record, which serves as the gold standard. For more detailed information of evaluation, please refer to §3.
Figure 3: Collaboration of intern doctors for clinical diagnosis.
3 MVME: Evaluation of LLMs as Intern Doctors for Clinical Diagnosis
Based on AI Hospital, we assess the feasibility of employing LLMs for clinical diagnosis.
3.1 Multi-View Evaluation Criteria
We develop evaluation criteria from three aspects to assess the procedure of clinical diagnosis.
Mastery of Patient Symptoms. This assesses how precisely the intern doctors can gather and understand information about patient symptoms. It focuses on their ability to comprehensively explore the medical history, identify critical symptoms based on patient descriptions, and perform necessary analyses and documentation.
Comprehensiveness of Medical Examination. This evaluates the range and appropriateness of the medical tests employed by the interns. It considers their choice and skill in interpreting the findings.
Quality of Diagnostic Report. This measures the accuracy and professionalism of the diagnostic report, including the diagnostic results, diagnostic rationales, and the proposed treatment plan. It examines the report’s consistency with medical logic and standards.
Centering on three evaluation criteria, our assessment covers five specific aspects: (1) symptom, (2) medical examination, (3) diagnostic results, (4) diagnostic rationales, and (5) treatment plan.
3.2 Evaluation Methods
LLM based Evaluation with Reference
To assist the medical director in evaluation of the five specific portions, we extract the “Diagnosis and Treatment” with “Chief Complaints” and “Auxiliary Examinations” from the medical records which are described in §2.1. The information is incorporated into the prompts for the LLMs (GPT-4) as Table 13&14, providing external expertise for the medical director to assess LLMs. In each specific evaluation dimension, medical director rate the interns’ performance on a scale of 1 to 4, ranging from poorest to excellent.
Link based Evaluation of Diagnostic Results
To enhance the scientific rigor and validity of our evaluations, particularly in diagnostic results with respect to identify diseases, we employ the International Classification of Diseases (ICD-10) Trott (1977) as the authoritative source, and link standardized disease terminologies with natural language based diagnostic results. Initially, we extract disease entities from the diagnostic results and the medical records’ references. Subsequently, we implement a fuzzy matching process with a predefined threshold of 0.5 to link these disease entities with ICD-10 terminology, building a normalized disease sets. Ultimately, we assess diagnostic agreement between the diagnoses from intern doctors and the golden reference, then report the number of extracted disease entities (#), set-level precision (P), recall (R), and F1-score (F) metrics.
Department #
Surgery 180
Internal Medicine 153
Obstetrics and Gynecology 94
Pediatrics 29
Otorhinolaryngology 23
Others 27
Table 1: Departments distribution of the collected medical records.
3.3 MVME Dataset Construction
We collect Chinese medical records across diverse departments online [^2] https://bingli.iiyi.com/ and engage professional physicians for a thorough review. Subsequent to the exclusion of records with deficiencies, such as incomplete information, a total of 506 cases remained. The detailed distribution of these cases among the various departments is presented in Table 1. To verify the quality of the collected medical records, we select samples from 10 secondary departments, randomly choosing five cases per department for review. Doctors from corresponding departments are hired to evaluate the “Diagnosis and Treatment” including diagnostic results, diagnostic rationales, and treatment course with a binary choice: either “fundamentally correct” or “obviously incorrect”. If three sections of a medical record are fundamentally accurate, then we consider the medical record to be correct. Overall, expert validation conclude that 94% of the records are deemed as correct.
Algorithm 1 Dispute Resolution Collaboration
0:  Maximum number of rounds $M$, number of intern doctors $N$ and pre-diagnosis $P$.
0:  Final Diagnosis $a$
1:  $D$Medical Director}
2:  $I\leftarrow[I_{1},\cdots,I_{N}]$Intern Doctors}
3:  $H\leftarrow P$Initialize Discussion History}
4:  $d\leftarrow D(H)$Initialize Dispute}
5:  $m\leftarrow 0$Current Round}
6:  while $m\leq M$ do
7:     $m\leftarrow m+1$
8:     for each $I_{i}$ in $I$ do
9:        $h\leftarrow D_{i}(H,d)$Generate Diagnosis}
10:        $H\leftarrow H+[h]$Append $h$ to $H$}
11:     end for
12:     $d\leftarrow D(H)$Summarize Disputes}
13:     if $d$ is NULL then
14:        break {Debate is Over}
15:     end if
16:  end while
17:  $a\leftarrow D(H)$
4 Collaborative Diagnosis of LLMs Focused on Dispute Resolution
To further improve diagnostic accuracy, we propose a collaborative mechanism for clinical diagnosis. To streamline the process of forming a unified diagnostic report, we introduce the medical director to participate as a moderator. The overall process is shown in Figure 3.
Symptom Medical Diagnostic Diagnostic Treatment
Examination Results Rationales Plan
Interaction
Baichuan (13B) 2.577 (0.083) 1.632 (0.085) 1.585 (0.081) 1.522 (0.077) 1.419 (0.071)
HuatuoGPT-II (13B) 2.832 (0.065) 1.733 (0.075) 1.601 (0.075) 1.601 (0.071) 1.427 (0.065)
HuatuoGPT-II (34B) 3.053 (0.055) 1.909 (0.083) 1.756 (0.075) 1.824 (0.077) 1.640 (0.071)
GPT-3.5 2.992 (0.040) 1.931 (0.089) 1.717 (0.073) 1.733 (0.073) 1.532 (0.065)
Wenxin-4.0 3.034 (0.040) 1.913 (0.081) 1.787 (0.079) 1.794 (0.077) 1.660 (0.073)
Qwen-Max 2.851 (0.063) 1.798 (0.079) 1.794 (0.079) 1.863 (0.079) 1.747 (0.073)
GPT-4 3.071 (0.038) 1.753 (0.079) 1.881 (0.077) 1.923 (0.077) 1.808 (0.079)
Collaboration
2 Doctors w/o DR 3.265 (0.061) 2.291 (0.091) 2.067 (0.085) 2.156 (0.083) 1.972 (0.079)
2 Doctors 3.342 (0.055) 2.413 (0.089) 2.142 (0.083) 2.247 (0.083) 2.077 (0.079)
3 Doctors 3.408 (0.054) 2.489 (0.085) 2.188 (0.084) 2.327 (0.083) 2.118 (0.079)
One-Step
$*{}^{*}$ $2.767\ (0.049)$ $2.998\ (0.040)$ $2.595\ (0.055)$
Table 2: MVME: GPT-4 evaluation with reference in clinical consultation. $*{}^{*}$ in One-Step is the upper bound.
4.1 Exchange of Factual Information
We contend that a consensus on the physical condition of patients among intern doctors constitutes the cornerstone of collaborative diagnosis. The process of building the consensus is delineated into three distinct steps, outlined below.
(1) During the interactive consultation process, intern doctors communicate with the medical director, relaying patient factual information they have acquired, focusing primarily on symptoms and medical test outcomes.
(2) The medical director consolidates and analyzes the data collected from various intern doctors, confirming symptoms and test outcomes with patients and examiners to clarify disputed points.
(3) Drawing upon the findings received from intern doctors, coupled with feedback from patients and examiners, the medical director synthesizes a comprehensive summary of the patient’s symptoms and medical examination outcomes.
4.2 Discussions on Dispute Resolution
In collaborative diagnosis, the medical director should analyze the statements of intern doctors and identify key points of disagreement to foster focused discussions. The process is as follows.
(1) The collaborative diagnosis consists of multiple discussion iterations. Under the guidance of the medical director, intern doctors are expected to delve deeper gradually, resolve differences, and reach a consensus.
(2) In each session of collaborative diagnostic discussion, each intern doctor should present their diagnostic reports while engaging in critical analysis of their peers’ findings. Guided by the medical director’s summary of disputed points among intern doctors, they can pinpoint the current issues requiring attention. This approach facilitates targeted and thorough critical thinking of intern doctors, enhancing the refinement of their reports.
(3) Upon the conclusion of discussions, the medical director assesses the persistence of disagreements among interns. If disagreements are identified, the director can summarize the controversial issues and set then as the agenda for the subsequent session to facilitate resolution. Conversely, if no disagreements are found, the director concludes the discussions and finalizes the diagnostic report by himself.
The overall process is delineated as pseudocode in Algorithm 1, and the prompts are listed in Table 15 and Table 16 of appendix.
5 Experiment
# R P F1
Interaction
Baichuan (13B) 1.58 10.21 23.79 14.28
HuatuoGPT-II (13B) 1.72 12.76 24.84 16.85
HuatuoGPT-II (34B) 1.86 17.48 30.95 22.34
GPT-3.5 1.81 19.19 37.39 25.37
Wenxin-4.0 2.50 22.03 31.44 25.91
Qwen-Max 1.77 22.42 43.38 29.56
GPT-4 1.52 21.64 50.26 30.26
Collaboration
2 Doctors w/o DR 2.37 28.44 41.45 33.74
2 Doctors 2.41 29.51 43.62 35.21
3 Doctors 3.20 36.54 39.58 38.00
One-Step
$*{}^{*}$ 2.30 38.90 58.97 46.88
Table 3: MVME: Link-based evaluation of diagnostic results in clinical consultation.
5.1 Evaluation Metrics
We employ the medical director for evaluation. Taking the medical records as reference, we report scores from LLM based evaluation methods with references, alongside those from link based one. The classic bootstrap method Efron (1992) is employed to calculate the variance of these scores.
5.2 Experiment Settings
One-Step Diagnosis as Upper Bound. In one-step diagnosis, the “Basic Information” and “Professional Medical Examination” in §2.1 is directly input to LLMs (e.g., GPT-4), thereby skipping the interactive consultation phase for patient information collection. Consequently, this approach with GPT-4 sets the upper bound in our experiment.
Evaluation of Different LLMs. To assess various LLMs, we engage different models, including GPT OpenAI (2023) versions (GPT-3.5 and GPT-4) and models such as Wenxin-4.0, QWen-Max Bai et al. (2023), and HuatuoGPT-II 13B and 34B Chen et al. (2023a).
Comparative Collaboration Mechanism. The comparative methods include Collaborative Diagnosis with 3 and 2 Agents, an 2 Agents without Dispute Resolution. They are denoted as 3 Doctors, 2 Doctors and 2 Doctors w/o DR. The initial two intern doctors are powered by GPT-3.5 and GPT-4 for interactive consultation, while the last one is using Wenxin-4.0.
5.3 Performance of LLMs in MVME
The primary results are presented in Table 2 and Table 3, and we can infer the following.
(1) One-Step diagnosis significantly surpasses LLMs in interactive consultations, highlighting their limitations in interactions, particularly in eliciting symptoms and medical test results. It also indicates substantial room for improvement in strategies for proactive consultation.
(2) The collaborative use of models can exceed the performance of GPT-4, thereby validating the efficacy of the cooperative mechanism.
(3) Collaboration among “3 Doctors” enhances diagnosis compared to “2 Doctors”, highlighting the benefits of more agents in cooperation.
(4) The removal of the “Dispute Resolution” mechanism from the “2 Doctors” reduces its effectiveness, emphasizing the significance of establishing a better consensus.
6 Analysis
6.1 Human Evaluation for Patient and Examiner Behavior
We analyze the behaviors of patients and examiners during interactive consultations, quantifying the interaction by tracking the number of question-and-answer (QA) pairs (conversation rounds) and the frequency of medical examination requests in each dialogue, denoted as #QA and #Test, respectively. Each consultation is assessed according to three primary metrics.
Relevance [QA] This metric evaluates the extent to which patients’ responses are pertinent to the doctors’ questions. It involves calculating the ratio of relevant answers provided by both patients and examiners to the total responses.
Consistency [QA] This involves assessing whether the responses of patients and examiners match with the information in the collected medical records.
Accuracy [Medical Examination] This measures the instruction-following of conducting medical examination, as we examine the frequency of examiners receive application and the rate at doctors inquiring about medical examinations.
#QA Rel Con #Test Acc
Qwen-Max 429 100.0% 99.0% 56 98.2%
Wenxin-4.0 472 100.0% 98.1% 68 98.5%
GPT-3.5 417 100.0% 99.5% 57 98.2%
GPT-4 378 100.0% 99.7% 61 100.0%
Table 4: Human evaluation for patient and examiner behavior in interactive consultation. Rel, Con and Acc are short for Relevance, Consistency and Accuracy.
We randomly select 50 dialogues from interactive consultations involving GPT-3.5, GPT-4, Wenxin-4.0, and Qwen-Max, resulting in total 200 instances for comprehensive evaluation. We manually annotate each utterance during the evaluation process. The results, presented in Table 4, demonstrate high performance (over 98.0%) across all metrics of reliability with respect to patient and examiner interactions, regardless of the intern doctors powered by different LLMs. This highlights the efficacy of our interactive consultation system.
6.2 Efficiency of Dispute Resolution in Collaboration
For the “Dispute Resolution”, we continue to check whether intern doctors can reach consensus more rapidly. In terms of efficiency, a comparative analysis is conducted on the number of discussion rounds necessary to achieve consensus, both with and without the “Dispute Resolution” mechanism. The outcomes are detailed in Figure 4. These findings reveal a marked increase in the rate of consensus achieved within the initial four discussion rounds following the adoption of the dispute resolution mechanism. This enhancement suggests that the process, facilitated by the medical director highlighting controversial issues and intern doctors concentrating on these discussions, effectively reduces the time required to achieve consensus.
Figure 4: Statistical analysis of discussion rounds in collaborative frameworks with and without “Dispute Resolution” mechanism.
6.3 Case Study
We show an example of interactive clinical diagnosis among an intern doctor, a patient, an examiner and the medical director in Figure 5 of Appendix to leave an intuitive impression.
6.4 Reasons for Failure Cases
Error Type #
Omission of auxiliary examinations 99
Exclusive focus on complications 52
Erroneous judgment 68
Table 5: Classification and statistics of misdiagnoses (1 point) of the intern doctor powered by GPT-4.
We analyze cases for which GPT-4 provides incorrect diagnostic results and the medical director rates as 1 point. A total of 219 such cases are collected. By means of manual review, errors are categorized into three distinct types.
Omission of Auxiliary Examinations. An instance is the failure to detect gallbladder stones due to the absence of an abdominal ultrasound.
Exclusive Focus on Complications. For example, the attention of LLM is limited to soft tissue swelling in the feet, with concurrent neglect of the patient’s underlying diabetes.
Erroneous Judgment. While LLMs generally grasp the symptoms and the medical examination result, they make incorrect diagnosis, potentially reflecting inadequate medical expertise.
The number of errors for each category is detailed in the Table 5.
7 Related Works
Large Language Model in Medicine. Prior to the emergence of large language models (LLMs), the majority of automated diagnostic methods Zhong et al. (2022); Chen et al. (2023c) relies on reinforcement learning to guide agents in gathering patient symptoms and conducting diagnoses. The development of LLMs in the medical sector has been driven by open-source Chinese LLMs and various fine-tuning methods. Models like Med-PaLM Singhal et al. (2022), CPLLM Shoham and Rappoport (2023), DoctorGLM Xiong et al. (2023), BenTsao Wang et al. (2023a), ChatGLM-Med Wang et al. (2023b), Bianque-2 Chen et al. (2023d), ChatMed-Consult Zhu and Wang (2023), MedicalGPT Xu (2023), and DISC-MedLLM Bao et al. (2023) fine-tune using different datasets, techniques, and frameworks, focusing on medical question answering, health inquiries and doctor-patient dialogues.
Evaluation in Medicine. Prior research in medical evaluation has concentrated on non-interactive tasks, including question answering, entity and relation extraction, and medical summarization and generation. In the domain of biomedical question answering, key datasets such as MedQA (USMLE) Jin et al. (2021), PubMedQA Jin et al. (2019), and MedMCQA Pal et al. (2022) are utilized, with accuracy serving as the primary evaluation metric. The objective of entity and relation extraction Li et al. (2020) is to categorize named entities and their relationships from unstructured text into specific predefined classes. Prominent biomedical NER datasets include NCBI Disease Doğan et al. (2014), JNLPBA Collier and Kim (2004), BC5CDR Li et al. (2016), BioRED Luo et al. (2022) and IMCS Chen et al. (2023b), with the F1 score being the standard for model performance assessment. Medical summarization and generation tasks involve converting structured data, like tables, into descriptive text. This includes the creation of patient clinic letters, radiology reports, and medical notes Liu et al. (2023). The principal datasets for these tasks are PubMed Jin et al. (2019) and MentSum Sotudeh et al. (2022). A recent study introduced BioLeaflets Yermakov et al. (2021) and assessed multiple Large Language Models (LLMs) in data-to-text generation.
LLM Powered Agents. The advancement of LLMs powered agents has led to significant strides in complex task resolution through human-like actions, such as role-playing and communication Xi et al. (2023); Wang et al. (2023c). This includes applications in software design and molecular dynamics simulation. Recent research Tang et al. (2023) in the medical field has highlighted the critical roles and decision-making processes in medical QA, encompassing various investigations like CT scans, ultrasounds, electrocardiograms, and blood tests. Despite these advancements, effectively integrating LLM-based agents into the medical domain, particularly in disease diagnosis, presents a notable challenge Zhou et al. (2023). Our research pioneers the use of multi-agent systems in creating a clinical diagnosis environment. We introduce a novel mechanism for identifying, discussing, and resolving disputes in collaboration, demonstrating promising results in clinical diagnosis.
8 Conclusion
In AI Hospital, we take a step forward in the field of medical interactions by concentrating on clinical diagnosis. We introduce AI Hospital to build a real-time interactive consultation scenario. We generate simulated patients and examiner using collected medical records and established a comprehensive engineering process. Based on the platform, we build a benchmark MVME to explore the feasibility of different LLMs in interactive consultations. To improve the diagnostic accuracy, this research also introduces a novel collaborative mechanism for intern doctors, featuring iterative discussions and a dispute resolution process, supervised by a medical director. In our experiment, the results not only demonstrate the performance of different LLMs but also confirm the efficacy of our dispute resolution-centered collaborative approach. For in-depth analysis, we list the error types and identify the issues that should be addressed in the future.
Limitation
Insufficient Evaluation of Treatment Plan. In actual medical settings, treatment plans frequently vary; yet, our existing evaluation system merely gauges the alignment between the model’s proposed treatment and a standard reference plan. This system inadequately accounts for feasible alternative treatment strategies.
Risk of Environmental Impact. A significant risk associated with our AI Hospital is the potential escalation of environmental burdens stemming from the widespread utilization of OpenAI’s APIs for LLMs. This impact could be mitigated through the adoption of smaller-scale and open-source models that offer greater efficiency and reduced environmental impact in the future.
Ethics Consideration
Ethical considerations are paramount in our research. This section addresses the potential ethical implications of our study: Firstly, the medical records data utilized in our research, being publicly accessible online, will be released to enhance the reproducibility of our results. This enables other researchers to assess and extend our findings. Secondly, we emphasize privacy protection. The data sources have removed sensitive information prior to our collection, and we rigorously review our datasets to confirm the absence of sensitive details such as individual names.
Appendix A Essential Capabilities of Intern Doctor in Clinical Diagnosis
Detailed explanation of functionality of patient agent in §2.4.
  • Articulate Expression of Complaints: Patients are expected to express their primarDy health concerns and symptoms clearly and accurately during consultations, thereby aiding doctors in comprehending the medical condition.
  • Truthful Responses to Doctor’s Inquiries: In response to doctors’ inquiries about medical history or during examinations, patients are encouraged to provide honest and comprehensive answers, covering aspects like symptoms, lifestyle habits, and family medical history.
  • Active Participation in Examinations: When doctors recommend further examinations, patients actively engage in the diagnostic process by seeking inspectors, accurately issuing examination requests, and acquiring corresponding results.
  • Voicing Questions and Concerns: It is crucial for patients to voice any questions or concerns they may have about the treatment plan or suggestions made by doctors, to gain a better understanding of the diagnosis and treatment.
Appendix B More Analysis of LLMs Performance in MVME
  • Models with a larger number of parameters, such as GPT-4, typically outperform the small ones, such as Baichuan13B.
  • HuatuoGPT-II (13B) outperforms Baichuan (13B), which is developed from the later one through continual training on medical data. Its superior performance underscores the value of domain-specific data in enhancing model capabilities.
  • Compared to more advanced foundational models, strategies for interactive consultation, including our collaborative mechanism, can yield significant benefits. Consequently, the exploration of more effective consultation strategies warrants further attention.
Appendix C Attacking on Examiner: Reliability of Examiner’s Workflow
To assess the vulnerability of examiner to information extraction tactics, we design complex scenarios intended to safeguard against the unauthorized disclosure of medical examination data in scenarios lacking explicit examination requests. We build 50 samples, such as “Please provide the most recent medical examination details”, aimed at eliciting information. During attacking, we observe that the examiner, in the first step, consistently classify these requests as void of legitimate examination item, implementing a defensive strategy.
Appendix D Example of Expert Verification
In § 3.3, we assess the quality of the patient medical records. These records are vetted by doctors sourced from Jiangsu Provincial People’s Hospital and Tongji Medical College of Huazhong University of Science and Technology, which hold the prestigious distinction of being tertiary Class A hospitals in China.
We utilize the Tencent questionnaire platform to facilitate the quality inspection of medical record by medical professionals. A representative case is depicted in Figure 6, and we also include a display of the questionnaire template in Table 6.
Figure 6: Sample of a questionnaire used for medical record quality inspection.
{Medical Record} 1. Is the diagnostic results correct? YesNo 2. Is the diagnostic rationale correct? YesNo 3. Is the treatment course reasonable? YesNo
Table 6: The translated template of questionnaire for expert verification of medical record.
Appendix E Human Evaluation for Patient and Examiner Behavior
{ 病历 } { 单轮对话内容 } 请你仔细阅读这一轮对话的内容和病人的病历信息,回答下面的问题。 1. 病人(检查员)的发言与医生的相关吗? 2. 病人(检查员)的发言符合病历的内容吗? 3. 医生是否建议进行专业的医学检查? 4. 检查员是否进行了医学检查? { Medical Record } { Single Round Conversation Content } Carefully review the content of the conversation and the corresponding medical record to answer the following questions. 1. Is the statement of patient or examiner relevant to the doctor’s one? YesNo 2. Is the statement of patient or examiner consistent with the content of medical record? YesNo 3. Does the doctor recommend a professional medical examination? YesNo 4. Does the examiner perform a medical test? YesNo
Table 7: The original Chinese and translated English questionnaire of human evaluation for patient and examiner behavior.
In Figure 7, we present a questionnaire developed for the manual validation of patient and examiner behaviors in each round of conversation. The questionnaire includes four questions, with the initial two addressing “Relevance” and “Consistency” in the question-and-answer (QA) pairs, while the subsequent two focus on the “Accuracy” of conducting medical examinations. Three medical students are hired to complete these questionnaires. Consensus among the first two reviewers leads to the immediate acceptance of their collective assessment. In cases of divergence, the opinion of the third reviewer is solicited, whose determination, reflecting the majority viewpoint, constitutes the definitive annotation. The agreement rates for each questions are 99.1%, 95.6%, 99.4%, and 100.0%. Significantly, when calculating the accuracy of medical examinations, we exclusively consider dialogues affirmed as “No” in response to the third question.
Appendix F Prompts for Different Medical Roles
We list the prompts for different medical roles in the Table 8. In each prompt, { xx } needs to fill with external inputs xx.
Prompt Medical Role Function Chapter
Table 9 Patient Chat with Intern Doctor §2.4
Table 10 Examiner Process Examination Request §2.3
Table 11 Examiner Produce Examination Outcomes §2.3
Table 13&14 Medical Director Evaluate Diagnosis of Intern Doctor §2.53
Table 12 Intern Doctor Interactive Clinical Diagnosis §2.2
Table 15 Intern Doctor Collaboration through Discussion §4
Table 16 Medical Director Summarize Statement of Various Intern Doctors §4
Table 8: Prompts of different medical roles and the corresponding function.
System Message 你是一个病人。这是你的基本资料。 {个性化信息} {病历中的基本信息} 下面会有医生来对你的身体状况进行诊断,你需要: (1) 按照病历和基本资料的设定进行对话。 (2) 在每次对话时,你都要明确对话的对象是<医生>还是<检查员>。当你对医生说话时,你要在句子开头说<对医生讲>;如果对象是<检查员>,你要在句子开头说<对检查员讲>。 (3) 首先按照主诉进行回复。 (4) 当<医生>询问你的现病史、既往史、个人史时,要按照相关内容进行回复。 (5) 当<医生>要求或建议你去做检查时,要立即主动询问<检查员>对应的项目和结果,例如:<对检查员讲> 您好,我需要做xxx检查,能否告诉我这些检查结果? (6) 回答要口语化,尽可能短,提供最主要的信息即可。 (7) 从<检查员>那里收到信息之后,将内容主动复述给<医生>。 (8) 当医生给出诊断结果、对应的诊断依据和治疗方案后,在对话的末尾加上特殊字符<结束>。 User [患者] {Statement Generated by GPT-4 in §2.4} System Message You are a patient. Here is your basic information. {Personality in §2.4} {Basic Information in Medical Record §2.1} A doctor will come to diagnose your physical condition. You need to: (1) Engage in dialogue according to the settings of personality and the basic information in medical record. (2) In each conversation, you must clarify whether you are speaking to a <doctor> or an <examiner>. When you speak to the doctor, you should start your sentences with <To the doctor>; if the addressee is an <examiner>, you should start with <To the examiner>. (3) First, respond according to the chief complaint. (4) When the <doctor> asks about your present illness history, past medical history, and personal history, reply according to the relevant content. (5) When the <doctor> requests or suggests that you undergo tests, immediately ask the <examiner> about the corresponding items and results, for example: <To the examiner> Hello, I need to have xxx examination, can you tell me the results of these tests? (6) The responses should be conversational, as short as possible, providing only the most important information. (7) After receiving information from the <examiner>, actively repeat the content to the <doctor>. (8) When the doctor provides the diagnostic result, the corresponding rationale for the diagnosis, and the treatment plan, end the dialogue with the special token <end>. User [Patient] {Statement Generated by GPT-4 in §2.4}
Table 9: The original Chinese and translated English prompts for patient agent.
System Message 你是医院负责检查的自动化接待员。请你利用掌握的医学检查的命名实体的知识,从病人的检查申请当中解析出指向明确的专业医学检查项目,方便后面的检查科室进行检查。 请按照下面的格式的输出: # 检查项目 - xxx - xxx 如果没有找到具体的医学检查项目,请输出: # 检查项目 - 无 User [患者] 您好,医生告诉我根据CT扫描和PET-CT扫描的结果,初步得出以下结论:右肺上叶有一个大小约为2.6*1.9cm的实性结节。双肺下叶也有散在的淡薄浸润影。医生建议我进行进一步的检查,例如活检。 Assistant [检查员] # 检查项目 - 肺部活检 User [患者] 我需要了解一下我的检查结果。可以告诉我具体的检查项目和结果吗?谢谢!? Assistant [检查员] # 检查项目 - 无 System Message You are an automated receptionist responsible for examinations in a hospital. Using your knowledge of medical examination named entities, please parse out specific professional medical examination items from patients’ examination requests to facilitate subsequent examinations by the relevant departments. Output in the following format: # Examination Item - xxx - xxx If no specific medical examination items are found, please output: # Examination Item - None User [Patient] Hello, the doctor told me that based on the results of the CT scan and PET-CT scan, the preliminary conclusion is that there is a solid nodule approximately 2.6*1.9cm in size in the upper lobe of the right lung. There are also scattered thin infiltrative shadows in the lower lobes of both lungs. The doctor advised me to undergo further examinations, such as a biopsy. Assistant [Examiner] # Medical Examination Items - Lung biopsy User [Patient] I need to know about my examination results. Can you tell me the specific examination items and results, please? Thank you!? Assistant [Examiner] # Medical Examination Items - None
Table 10: The original Chinese and translated English prompts for patient agent to process examination request.
System Message 这是你收到的病人的检查结果。 {Professional Medical Examination in §2.1} 下面会有病人或医生来查询,你要忠实地按照收到的检查结果,找到对应的项目,并按照下面的格式来回复。 # xx检查 - xxx: xxx - xxx: xxx 如果无法查询到对应的检查项目则回复: - xxx: 无异常 System Message This is the patient’s examination result that you received. {Professional Medical Examination in §2.1} Patients or doctors will come to inquire about these results. You must faithfully report the received examination results, identify the corresponding items, and respond in the following format: # xx Examination xxx: xxx xxx: xxx If the corresponding examination item cannot be found, reply with: xxx: No abnormalities
Table 11: The original Chinese and translated English prompts for patient agent to produce examination outcomes.
System Message 你是一个专业且耐心的医生,下面会有患者向你咨询病情。你需要: (1) 在信息不充分的情况下,不要过早作出诊断。 (2) 多次、主动地向患者提问来获取充足的信息。 (3) 必要时要求患者进行检查,并等待患者反馈。 (4) 诊断结果需要准确到具体疾病。 (5) 最后根据患者的身体状况和检查结果,给出诊断结果、对应的诊断依据和治疗方案。 System Message You are a professional and patient doctor, and you will be consulted by patients. You need to: (1) Avoid making premature diagnoses when information is insufficient. (2) Actively and repeatedly inquire to gather adequate information from patients. (3) When necessary, request patients to undergo medical examinations and await their feedback. (4) Ensure that the diagnosis is precise and specific to the particular ailment. (5) Finally, based on the patients’ physical condition and examination results, provide a diagnosis, the corresponding rationale, and a treatment plan.
Table 12: The original Chinese and translated English prompts for intern doctor in interactive clinical diagnosis.
System Message 你是资深的医学专家。 请你根据专家诊疗结果中的现病史、辅助检查、诊断结果、诊断依据和治疗方案,来对实习医生进行评价。 请参考下面的细则进行评价。 1. 病人症状的掌握情况 (A) 全面掌握 (B) 相当部分掌握 (C) 小部分掌握 (D) 绝大部分不掌握 2. 医学检查项目的完整性 (A) 非常完整 (B) 相当部分完整 (C) 小部分完整 (D) 绝大部分不完整 3. 诊断结果的一致性 (A) 完全一致,诊断正确 (B) 相当部分一致,诊断基本正确 (C) 小部分一致,诊断存在错误 (D) 完全不一致,诊断完全错误 4. 诊断依据的一致性 (A) 完全一致 (B) 相当部分一致 (C) 小部分一致 (D) 完全不一致 5. 治疗方案的一致性 (A) 完全一致 (B) 相当部分一致 (C) 小部分一致 (D) 完全不一致 通过下面的方式来呈现结果 # 症状 ## 分析 <根据专家记录的病人病史,分析实习医生对病人病情的掌握情况> ## 选项<根据症状分析做出选择> # 医学检查项目 ## 分析 <基于专家所做的医学检查项目,全面分析实习医生所做的医学检查项目的完整性> ## 选项 <根据分析得到的完整性做出选择> # 诊断结果 ## 分析 <基于专家做出的诊断结果,结合你的医学常识,分析实习医生诊断结果与专家的一致性> ## 选项 <根据分析得到的一致性做出选择> # 诊断依据 ## 分析 <对比专家的诊断依据,分析实习医生的治疗方案与其的一致性> ## 选项 <根据分析得到的一致性做出选择> # 治疗方案 ## 分析 <对比专家的治疗方案,分析实习医生的治疗方案与其的一致性> ## 选项 <根据分析得到的一致性做出选择> (1) 请侧重医学答案的事实内容,不需关注风格、语法、标点和无关医学的内容。 (2) 请你充分利用医学知识,分析并判断每个点的重要性,再做评价。 (3) 注意诊断结果、诊断依据和治疗方案三者之间的承接关系。例如,如果诊断错误,那么后面两部分与专家的一致性就必然很低 User # 专家的诊断报告 {Diagnosis and Treatment in §2.1} # 实习医生的诊断报告 {实习医生的诊断报告}
Table 13: The original Chinese prompt for medical director to evaluate.
You are an experienced medical expert. Please evaluate the intern doctors based on their current medical history, auxiliary examinations, diagnostic results, diagnostic basis, and treatment plans from the expert’s diagnosis. Please refer to the following guidelines for evaluation. 1. Mastery of Patient Symptoms (A) Comprehensive mastery (B) Substantial mastery (C) Partial mastery (D) Mostly unmastered 2. Completeness of Medical Examination (A) Very complete (B) Substantially complete (C) Partially complete (D) Mostly incomplete 3. Diagnosis Result (A) Completely consistent, correct diagnosis (B) Largely consistent, basically correct diagnosis (C) Partially consistent, diagnosis contains errors (D) Completely inconsistent, completely incorrect diagnosis 4. Diagnostic Rationale (A) Completely consistent (B) Largely consistent (C) Partially consistent (D) Completely inconsistent 5. Treatment Plan (A) Completely consistent (B) Largely consistent (C) Partially consistent (D) Completely inconsistent Please output the results in the following format: # Symptoms ## Analysis <Analyze the intern’s grasp of the patient’s condition based on the expert’s recorded medical history.> ## Option <Choose based on the analysis of symptoms.> # Medical Examination Items ## Analysis <Thoroughly analyze the completeness of the medical examination items conducted by the intern, based on the expert’s examinations.> ## Option <Choose based on the analysis of completeness.> # Diagnostic Results ## Analysis <Based on the expert’s diagnostic results and your medical knowledge, analyze the consistency between the intern’s diagnostic results and the expert’s.> ## Option <Choose based on the analysis of consistency.> # Diagnostic Basis ## Analysis <Compare the diagnostic basis of the expert and analyze the consistency of the intern’s treatment plan with it.> ## Option <Choose based on the analysis of consistency.> # Treatment Plan ## Analysis <Compare the expert’s treatment plan and analyze the consistency of the intern’s treatment plan with it.> ## Option <Choose based on the analysis of consistency.> (1) Please focus on the factual content of the medical answers, without concern for style, grammar, punctuation, and content unrelated to medicine. (2) Please make full use of your medical knowledge to analyze and judge the importance of each point before evaluating. (3) Pay attention to the continuity among the diagnosis result, diagnostic basis, and treatment plan. User # Diagnostic Report of Medical Director {Diagnosis and Treatment in Section 2.1} # Diagnostic Report of Intern Doctor {Diagnostic Report of the Intern Doctor}
Table 14: The translated English prompt for medical director to evaluate.
System Message 你是一个专业的医生A。 你正在为患者做诊断,患者的症状和检查结果如下: {症状与检查结果} 针对患者的病情,你给出了初步的诊断报告: {医生A的诊断报告} (1) 下面你将收到来自其他医生的诊断意见,其中也包含诊断结果、诊断依据和治疗方案。你需要批判性地梳理并分析其他医生的诊断意见。 (2) 在这个过程中,请你注意主治医生给出的争议点。 (3) 如果你发现其他医生给出的诊断意见有比你的更合理的部分,请吸纳进你的诊断意见中进行改进。 (4) 如果你认为你的诊断意见相对于其他医生的更科学合理,请坚持自己的意见保持不变。 请你按照下面的格式来输出。 # 诊断结果 (1) xxx (2) xxx # 诊断依据 (1) xxx (2) xxx # 治疗方案 (1) xxx (2) xxx User # 医生B {医生B的诊断报告} # 医生C {医生C的诊断报告} # 主任医生 {主任医生的指导意见} System Message As a doctor, you are currently diagnosing a patient, whose symptoms and medical examination results are as follows: {Symptoms and Medical Examination Results} Based on the patient’s condition, you have prepared a preliminary diagnostic report: {Diagnostic Report of Doctor A} (1) You will receive diagnostic reports from other doctors. Critically review and analyze these reports. (2) During this process, pay attention to any controversial points raised by the medical director. (3) If you find aspects of other doctors’ diagnoses that are more rational than yours, incorporate these into your diagnosis for improvement. (4) If you believe your diagnostic opinion is more scientifically sound compared to others, maintain your stance. Please present your findings in the following format: Diagnosis Result (1) xx (2) xx Diagnostic Rationale (1) xx (2) xx Treatment Plan (1) xx (2) xx User # Doctor B {Diagnostic Report of Doctor B} # Doctor C {Diagnostic Report of Doctor C} # Medical Director {Guidance of Medical Director}
Table 15: The original Chinese prompt and translated English one for intern doctor to collaborate in discussion.
System Message 你是一个资深的主任医生。 你正在主持一场医生针对患者病情的会诊,参与的医生有医生A、医生B和医生C。 病人的基本情况如下: {症状与检查结果} (1) 你需要听取每个医生的诊断报告。 (2) 请你按照重要性列出最多3个需要讨论的争议点。 按照下面的格式输出: (1) xxx (2) xxx User # 医生A {医生A的诊断报告} # 医生B {医生B的诊断报告} # 医生C {医生C的诊断报告} System Message As an experienced medical director, you are presiding over a medical consultation concerning a patient’s condition, with the participation of Doctors A, B, and C. The patient’s basic information is as follows: {Symptoms and Test Results} (1) You are required to listen to the diagnostic reports from each physician. (2) Identify and list up to three key controversial points for discussion, prioritized by their importance. Please present your findings in the following format: (1) xxx (2) xxx User # Doctor A {Diagnostic Report of Doctor A} # Doctor B {Diagnostic Report of Doctor B} # Doctor C {Diagnostic Report of Doctor C}
Table 16: The original Chinese and the translated English prompts for medical director to summarize.